OpenStudy (anonymous):

How to simplify algebraic ratio: (x^2-x-20)/(x^2+8x+16) I know how to simplify the smaller things like (x^2 - 16)/(x^2 - 4) but when it comes to the trinomials I'm lost.

hartnn (hartnn):

first, you need to factor the 2 trinomials, can you ?

OpenStudy (anonymous):

I believe they factor like: (x^2-x-20) = (x - 5) (x + 4) (x^2+8x+16) = (x + 4) (x +4)

hartnn (hartnn):

ohh! very good! thats correct :)

hartnn (hartnn):

so, what cancels out in numerator and denominator ?

OpenStudy (anonymous):

(x +4) and (x +4) So I'm left with (x +4) and (x - 5)

hartnn (hartnn):

yeah, (x-5) in numerator and x+4 in denominator, so your answer will be (x-5)/(x+4) thats it !

OpenStudy (anonymous):

Oh wow, that was easier than I thought! It looked difficult written in the text. Thank you!

hartnn (hartnn):

welcome ^_^